# SLTrain: a sparse plus low-rank approach for parameter and memory efficient pretraining

### v1

@@ 1 Introduction
Large foundation models have achieved tremendous success in various domains, including linguistics, computer vision and biology. In particular, large language models (LLMs), such as the GPT series [39, 5] and the LLaMA family [51, 52] have reshaped the perception of how machine understands human languages. The predominant success of these models is primarily due to the model size, usually scaling to hundreds of billions of parameters. The scaling laws seem to suggest the capacity of LLMs grows with the model size [25], but nonetheless requiring massive amount of resources for pre-training, storing, and fine-tuning. Particularly, memory requirement for training an LLM imposes a hard barrier for model deployment on commercial GPUs. For example, the LLaMA 7B model requires a memory cost of approximately 84G under 32-bit floating point, including 28G of parameter state and 56G of optimizer state for momentum-based optimizers, like Adam [59, 28].
Building a LLM (from scratch) for downstream tasks typically involves two phases, i.e., pretraining and fine-tuning. The goal of pretraining is to capture general language patterns and semantics, enabling the model to acquire useful representations of words and sentences. Common pretraining objectives include masked language modeling [26], next token prediction [39, 40], etc. Fine-tuning then tailors the learned model representations from pretraining to downstream tasks, adjusting its weights to enhance performance on specific objectives. Pioneered by LoRA [21], recent works have popularized low-rank finetuning of a given pretrained model ($W_{0}$), where $W_{0}$ is generally full-rank (i.e., pretrained without any constraints).
@@
The premise is that LLMs usually adapt to downstream tasks in a low-dimensional subspace, which allows to parameterize the update by low-rank factors. Low-rank finetuning requires minimal trainable parameters and significant reduction of memory and computation resources [10, 12]. A number of works [14, 53, 18, 29, 30, 34, 3] have emerged to further improve the efficiency and adaptation capacity of LoRA.
While most of the works have focused on exploiting low-rank structure for fine-tuning, only a few [27, 24, 43, 47] have considered pretraining with low-rank parameterized weights. It has been observed that the performance of low-rank training often lags behind full-rank training despite the great potential for improving training and memory efficiency [47, 59]. This is because neural networks often exhibit full-rank structure in the weights and imposing low-rank restrictions could significantly limit their representation power. Hence, recent works have explored full-rank training with low-rank updates. For instance, ReLoRA [32] periodically restarts LoRA, where the low-rank updates are merged with the weights from the last period. However, ReLoRA also requires a warm-start full-rank training to achieve competitive performance [32]. GaLore [59] takes a different route by enforcing a low-rank structure not on the weights but on the gradients. This allows the Adam optimizer states to be stored in a low-dimensional space. While being memory efficient, GaLore is not parameter efficient because it still performs full parameter update with “projected-back” low-rank gradients.
Parameter efficiency is a desirable property post-pretraining for model deployment, fine-tuning, and model storage. On the other hand, memory efficiency is necessary for training models with lower hardware requirements. While low-rank models achieve both parameter and memory efficiency, as discussed earlier, they do not perform well in general [47, 59]. Therefore, a natural question is: how can we adapt low-rank training to achieve comparable performance as full-rank training while maintaining both parameter and memory efficiency?
@@ 2 Background on low-rank pretraining
where $P_{t}$ is a projection matrix constructed by taking the largest left singular vectors of $G_{t}$. To reduce computational cost, $P_{t}$ is computed every several iterations and is stored in the middle. Although being memory efficient (as $M_{t}$ and $V_{t}$ are computed in the smaller dimension), GaLore is not parameter efficient (due to computation of $P_{t}M_{t}$ for updating $W_{t}$).
@@ 3.1 Motivation for sparse plus low-rank parameterization
Low-rank and sparsity are parsimonious modeling strategies for exploring low-dimensional weight matrices. The low-rank component aims to learn the low-dimensional bases or eigenspaces of the weights. The sparse component, on the other hand, identifies effective neuron-wise interactions and disregards non-expressive ones. In linear algebra terms, the low-rank component enforces sparsity of singular values, whereas the sparse component enforces sparsity of individual entries. In general, low-rank matrices are not sparse, and sparse matrices are not necessarily low-rank [6]. These concepts provide complementary information that should be explored simultaneously.
To further motivate the sparse plus low-rank modeling, in Figure 2, we illustrate different statistics from weight matrices of a pretrained LLaMA 60M model on C4 dataset (introduced later in Section 5). In Figure 2(a), we plot the singular values of weight matrices of different layers. The plot exhibits a fast decay of the singular values followed by a more stable decay of (smaller) singular values. This suggests that the top subspaces can be effective in model compression, and therefore, builds a case for low-rank modeling. However, the tail singular value distribution shows that low-rank modeling purely may not be sufficient. In order to better understand the tail part, in Figure 2(b), we visualize the magnitude of the attention output weight matrix before and after we extract the top $r$-dimensional subspaces ($r=128$) for the last attention layer. It is apparent that, after removing the top subspaces, both the magnitudes and the variation of the entries present in the residual matrix become smaller. Plotting the magnitudes of the entries in Figure 2(c) we see that 97% of the entries have a magnitude below 0.04. In Appendix A, we provide such visualizations to corroborate the findings for other layers as well. Overall, Figure 2 suggests that a sparse matrix with random support can approximate the residual well given the magnitudes do not vary too much across the entries.
@@
In Table 1, we perform an ablation study that verifies the feasibility of using a random sparse support for approximating the residual matrix. Specifically, we take $L_{0}$ as the best rank-$r$ approximation ($r=128$) for the pretrained weight matrix $W_{0}$ and evaluate the perplexity score (PPL) on the validation set.
@@
We see that compared to the full-rank pretrained model, low-rank approximation $L_{0}$ suffers from a drastic performance drop. We also augment the low-rank approximation $L_{0}$ with either top 3% or random 3% of entries of the residual matrix, which we label as top sparse or random sparse pruning, respectively. We observe that $L_{0}$ plus top sparse pruning performs better compared to $L_{0}$ plus random sparse pruning. Nonetheless, both the performance is poor. We further evaluate fixing the low-rank approximation (to $L_{0})$ and only optimizing the sparse components with either top support or random support (both run for five times). Averaged PPL (over the five runs) for both the approaches improve and are comparable. This shows that fixing a random support for the sparse factor is a promising strategy from both efficiency and performance point of view. We explore learning both the sparse and low-rank factors in the next section.
@@ 3.2 Our proposed modeling
where $B\in{\mathbb{R}}^{d\times r},A\in{\mathbb{R}}^{r\times p}$ are low-rank factors with $r<\min\{d,p\}$ being the rank parameter and $S\in{\mathbb{R}}^{m\times n}$ is a sparse matrix. The number of non-zero entries (${\mathrm{nnz}}$) in $S$ is determined by the sparsity level parameter $\delta\in(0,1)$, i.e., ${\mathrm{nnz}}(S)=\delta dp$. So, the total number of parameters for the proposed parameterization is $(d+p)r+\delta dp$, which is much smaller than the full-rank layer parameters $dp$ when we choose $\delta\ll 1$. In addition to being parameter efficient, the optimization states also cost less memory and scales linearly with the number of trainable parameters. Finally, we note that the overall rank of $W$ will generally be high due to the presence of the sparse factor $S$.
The performance of such a parameterization highly depends on whether there exists an implementation that is both computation and memory efficient. Nevertheless, modern GPU hardware is not suited for sparse tensor multiplication $Sx$ for given input $x$, as well as its gradient, especially when $S$ presents an unstructured sparsity pattern [7]. This causes increased computational bottleneck despite showing memory advantage. Thus, existing works on sparse network and training mostly rely on learning and storing a parameter mask (i.e., support) [48, 15, 33] by letting $S=M\odot U$, where $M\in\{0,1\}^{d\times p}$ is a binary mask and $U\in{\mathbb{R}}^{d\times p}$ is a dense parameter. This allows to exploit GPU accelerator for dense matrix computation. However, masking requires to learn and store both the support and a dense parameter for training, which significantly increases the memory cost.
In this work, we achieve memory efficiency by representing $S$ in terms of its indices and values, i.e., $({\mathcal{I}},{\mathcal{V}})\in{\mathbb{R}}^{2{\mathrm{nnz}}(S)}$. This is possible because we randomly (and uniformly) fix the support a priori. The motivation for using a random (but fixed) support comes from the usefulness of random support in Table 1. This ensures the memory scales with only the sparsity in $S$ (i.e., the support size) rather than the full size of $S$. The forward pass involves computing
@@ 3.3 Practical considerations
Initialization and scaling. We consider LoRA type of initialization for low-rank factors, i.e., Kaiming initialization [19] for $A$ factor and zero initialization for $B$ factor. For sparse factor, we adopt uniform initialization for the values ${\mathcal{V}}$ in the range of $[-1/\sqrt{d_{\rm in}},1/\sqrt{d_{\rm in}}]$, where $d_{\rm in}$ denotes input feature dimension. We choose the sparse support ${\mathcal{I}}$ uniformly at random up to the desired sparsity level $\delta$. Furthermore, in order to balance the contribution of low-rank factor and sparse factor, we follow LoRA [21] to scale the low-rank factors by $\alpha/r$ where the balancing parameter $\alpha$ is a hyperparameter.
@@
This hyperparameter, along with the stepsize has a joint effect on the training speed of low-rank versus sparse factors.
@@ 5.1 Pretraining LLMs
Memory cost estimation. We compare the proposed SLTrain with the low-rank baseline models in terms of estimated memory consumption. Following [59], we compute memory estimates with bfloat16 format, where each floating point number occupies 2 bytes. We remark that SLTrain stores the indices with int64 format, which occupies 8 bytes per number. The memory cost for a training algorithm consists of the parameter memory and optimizer state memory. The parameter memory refers to the memory occupied by storing parameters, and the optimizer state memory refers to the memory required to store the first and second-order moment statistics, e.g., in Adam. Table 2 reports the total estimated memory cost for each method. The detailed breakdown of memory estimation can be found in Appendix C.
Perplexity vs efficiency. In Table 2, we compare the performance of different methods in three aspects: perplexity score, parameter size, and memory cost. We observe that SLTrain performs comparatively as the full-rank training and GaLore [59] while achieving further reduction in parameter size and memory cost. In addition, SLTrain only adds a small parameter and memory overhead to the low-rank parameterization, yet significantly improves the perplexity score. Hence, learning the additional sparse factor indeed helps in strengthening the representation capacity of SLTrain. This intuition is also validated in Figure 9 (Appendix B), where we plot the singular value distribution for different weight matrices. Due to the presence of the sparse factor, we observe that the spectrum of the SLTrain weight matrices gets enhanced beyond $r=128$.
@@
Scaling to LLaMA 7B model. For pretraining the LlaMA 7B model, due to the resource constraints, we only compare SLTrain with GaLore, implemented with 8-bit Adam [9] on 4$\times$80G A100 GPUs, without per-layer weight updates nor gradient checkpointing. [^5] For full-rank model, 8-bit Adam throws out-of-memory error on 4$\times$80G A100 GPUs.. We directly use the training scripts of GaLore, available at https://github.com/jiaweizzhao/GaLore. In Table 4, we observe that SLTrain performs comparably to GaLore in terms of perplexity and throughput while achieving significant memory reduction of 26% per GPU device.
@@
How do rank $r$ and sparsity $\delta$ affect performance? In Table 5, we validate the performance of training the Llama 60M and 130M models by varying the hyperparameters $r$ and $\delta$. We notice that, in general, when more parameters are added (corresponding to higher $r$ and $\delta$), the performance is better. However, this performance gain is also accompanied by an increase in memory footprint.
Table 5: Ablation comparison with low-rank and sparse parameterization along with change of . Validation perplexity ($\downarrow$) and parameter size and estimated memory cost in brackets.
@@
5.2 Fine-tuning LLMs
In addition to pretraining, the sparse plus low-rank factor may also be used for fine-tuning LLMs as: $W=W_{0}+BA+S$, where $W_{0}$ is a given pretrained model weight, $B$ and $A$ are low-rank factors, and $S$ is the sparse factor. As in Section 3.2, we learn the fine-tuned weights $B,A,$ and $S$ and term this as SLTrain-FT. The fine-tuning experiments are conducted for RoBERTa base model (with 125M parameters) on GLUE benchmarks. We use $r=8$ for all methods and tune the hyperparameters of SLTrain-FT (discussed in Appendix D). The results on benchmark datasets are shown in Table 6. We observe that SLTrain-FT performs competitively as the baselines. While there is no specific memory advantage of SLTrain-FT when a general full-rank $W_{0}$ is given, memory reductions can be obtained when $W_{0}$ is learned via SLTrain.
CoLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg
Full-rank FT 62.24 90.92 91.30 79.42 94.57 87.18 92.33 92.28 86.28
LoRA (rank=8) 61.83 90.80 91.90 79.06 93.46 86.94 92.25 91.22 85.93

### v2

@@ 1 Introduction
Large foundation models have achieved tremendous success in various domains, including linguistics, computer vision and biology. In particular, large language models (LLMs), such as the GPT series [39, 5] and the LLaMA family [51, 52] have reshaped the perception of how machine understands human languages. The predominant success of these models is primarily due to the model size, usually scaling to hundreds of billions of parameters. The scaling laws seem to suggest the capacity of LLMs grows with the model size [25], but nonetheless requiring massive amount of resources for pre-training, storing, and fine-tuning. Particularly, memory requirement for training an LLM imposes a hard barrier for model deployment on commercial GPUs. For example, the LLaMA 7B model requires a minimum memory cost of approximately 42G under 16-bit floating point, including 14G of parameter state and 28G of optimizer state for momentum-based optimizers, like Adam [59, 28].
Building an LLM (from scratch) for downstream tasks typically involves two phases, i.e., pretraining and fine-tuning. The goal of pretraining is to capture general language patterns and semantics, enabling the model to acquire useful representations of words and sentences. Common pretraining objectives include masked language modeling [26], next token prediction [39, 40], etc.
@@
Fine-tuning then tailors the learned model representations from pretraining to downstream tasks, adjusting its weights to enhance performance on specific objectives. Pioneered by LoRA [21], recent works have popularized low-rank finetuning of a given pretrained model ($W_{0}$), where $W_{0}$ is generally full-rank (i.e., pretrained without any constraints). The premise is that LLMs usually adapt to downstream tasks in a low-dimensional subspace, which allows to parameterize the update by low-rank factors. Low-rank finetuning requires minimal trainable parameters and significant reduction of memory and computation resources [10, 12]. A number of works [14, 53, 18, 29, 30, 34, 3] have emerged to further improve the efficiency and adaptation capacity of LoRA.
While most of the works have focused on exploiting low-rank structure for fine-tuning, only a few [27, 24, 43, 47] have considered pretraining with low-rank weights. It has been observed that the performance of low-rank training often lags behind full-rank training despite the great potential for improving training and memory efficiency [47, 59]. This is because neural networks often exhibit full-rank structure in the weights and imposing low-rank restrictions could significantly limit their representation power. Hence, recent works have explored full-rank training with low-rank updates. For instance, ReLoRA [32] periodically restarts LoRA, where the low-rank updates are merged with the weights from the last period. However, ReLoRA also requires a warm-start full-rank training to achieve competitive performance [32]. GaLore [59] takes a different route by enforcing a low-rank structure not on the weights but on the gradients. This allows the Adam optimizer states to be stored in a low-dimensional space. While being memory efficient, GaLore is not parameter efficient because it still performs full parameter update with “projected-back” low-rank gradients.
Parameter efficiency is a desirable property post-pretraining for model deployment, fine-tuning, and model storage. On the other hand, memory efficiency is necessary for training models with lower hardware requirements. Despite the importance of both parameter and memory efficiency, these two goals are often pursued independently. While low-rank models achieve both parameter and memory efficiency, as discussed earlier, they do not perform well in general [47, 59]. Therefore, a natural question is:
how can we adapt low-rank training to achieve comparable performance as full-rank training while maintaining both parameter and memory efficiency?
@@ 2 Background on low-rank pretraining
where $P_{t}$ is a projection matrix constructed by taking the largest left singular vectors of $G_{t}$. To reduce computational cost, $P_{t}$ is computed every several iterations and is stored in the middle. Although being memory efficient (as $M_{t}$ and $V_{t}$ are computed in the smaller dimension), GaLore is not parameter efficient due to computation of $P_{t}M_{t}$ for updating $W_{t}$.
@@ 3.1 Motivation for sparse plus low-rank parameterization
Both low-rank and sparsity are parsimonious modeling strategies for exploring low-dimensional weight matrices. The low-rank component aims to learn the low-dimensional bases or eigenspaces of the weights. The sparse component, on the other hand, identifies effective neuron-wise interactions and disregards non-expressive ones. In linear algebra terms, the low-rank component enforces sparsity of singular values, whereas the sparse component enforces sparsity of individual entries. In general, low-rank matrices are not sparse, and sparse matrices are not necessarily low-rank [6]. These concepts provide complementary information that should be explored simultaneously.
Despite that low-rank modeling alone can have limited expressivity due to the low-rank structure it imposes, we show in the below proposition that low-rank plus a uniform sparse matrix with only $\Omega(\log n/n)$ number of entries is full-rank with high probability.
Proposition 1.
Consider a matrix $S\in\mathbb{R}^{n\times n}$ with support $\mathcal{S}$ sampled uniformly at random with probability $\delta\in(0,1)$, i.e., $\mathbb{P}[(i,j)\in\mathcal{S}]=\delta$, for all $i,j\in[n]$. Suppose $\delta=\Omega(\log n/n)$, then with probability at least $1-O(1/n)$, $BA+S$ is full rank for arbitrary randomly generated $B,A$.
To further motivate the sparse plus low-rank modeling, in Figure 2, we illustrate different statistics from weight matrices of a pretrained LLaMA 60M model on C4 dataset (introduced later in Section 5). In Figure 2(a), we plot the singular values of weight matrices of different layers. The plot exhibits a fast decay of the singular values followed by a more stable decay of (smaller) singular values. This suggests that the top subspaces can be effective in model compression, and therefore, builds a case for low-rank modeling. However, the tail singular value distribution shows that low-rank modeling purely may not be sufficient. In order to better understand the tail part, in Figure 2(b), we visualize the magnitude of the attention output weight matrix before and after we extract the top $r$-dimensional subspaces ($r=128$) for the last attention layer. It is apparent that, after removing the top subspaces, both the magnitudes and the variation of the entries present in the residual matrix become smaller. Plotting the magnitudes of the entries in Figure 2(c) we see that 97% of the entries have a magnitude below 0.04. In Appendix B and C, we provide such visualizations for other layers of LLaMA 60M and Pythia 70M to further corroborate the findings. Overall, the figures suggest that a sparse matrix with random support can approximate the residual well given the magnitudes do not vary too much across the entries.
@@
In Table 1, we perform an ablation study that verifies the feasibility of using a random sparse support for approximating the residual matrix. Specifically, we take $L_{0}$ as the best rank-$r$ approximation ($r=128$) for the pretrained weight matrix $W_{0}$ and evaluate the perplexity score (PPL) on the validation set. We see that compared to the full-rank pretrained model, low-rank approximation $L_{0}$ suffers from a drastic performance drop. We also augment the low-rank approximation $L_{0}$ with either top 3% or random 3% of entries of the residual matrix, which we label as top sparse or random sparse pruning, respectively.
@@
We observe that $L_{0}$ plus top sparse pruning performs better compared to $L_{0}$ plus random sparse pruning. Nonetheless, both the performance is poor. We further evaluate fixing the low-rank approximation (to $L_{0})$ and only optimizing the sparse components with either top support or random support (both run for five times). Averaged PPL (over the five runs) for both the approaches improve and are comparable. This shows that fixing a random support for the sparse factor is a promising strategy from both efficiency and performance point of view. We explore learning both the sparse and low-rank factors in the next section.
@@ 3.2 Our proposed modeling
where $B\in{\mathbb{R}}^{d\times r},A\in{\mathbb{R}}^{r\times p}$ are low-rank factors with $r<\min\{d,p\}$ being the rank parameter and $S\in{\mathbb{R}}^{m\times n}$ is a sparse matrix. The number of non-zero entries (${\mathrm{nnz}}$) in $S$ is determined by the sparsity level parameter $\delta\in(0,1)$, i.e., ${\mathrm{nnz}}(S)=\delta dp$. So, the total number of parameters for the proposed parameterization is $(d+p)r+\delta dp$, which is much smaller than the full-rank layer parameters $dp$ when we choose $\delta\ll 1$. In addition to being parameter efficient, the optimization states also cost less memory and scales with the number of trainable parameters. Finally, we note that the overall rank of $W$ will generally be high due to the presence of the sparse factor $S$, based on Proposition 1.
The performance of such a parameterization highly depends on whether there exists an implementation that is both computation and memory efficient. Nevertheless, modern GPU hardware is not suited for sparse tensor multiplication $Sx$ for given input $x$, as well as its gradient, especially when $S$ presents an unstructured sparsity pattern [7]. This causes increased computational bottleneck despite showing memory advantage. Thus, existing works on sparse network and training mostly rely on learning and storing a parameter mask (i.e., support) [48, 15, 33] by letting $S=M\odot U$, where $M\in\{0,1\}^{d\times p}$ is a binary mask and $U\in{\mathbb{R}}^{d\times p}$ is a dense parameter. This allows to exploit GPU accelerator for dense matrix computation. However, masking requires to store both the support and a dense parameter for training, which significantly increases the memory cost.
In this work, we achieve memory efficiency by representing $S$ in terms of its indices and values, i.e., $({\mathcal{I}},{\mathcal{V}})\in{\mathbb{R}}^{2{\mathrm{nnz}}(S)}$. This is possible because we randomly (and uniformly) fix the support a priori. The motivation for using a random (but fixed) support comes from the usefulness of random support in Table 1. This ensures the memory scales with only the sparsity in $S$ (i.e., the support size) rather than the full size of $S$. Further, the forward pass involves computing
@@ 3.3 Practical considerations
Initialization and scaling. We consider LoRA type of initialization for low-rank factors, i.e., Kaiming initialization [19] for $A$ factor and zero initialization for $B$ factor.
@@
For sparse factor, we adopt uniform initialization for the values ${\mathcal{V}}$ in the range of $[-1/\sqrt{d_{\rm in}},1/\sqrt{d_{\rm in}}]$, where $d_{\rm in}$ denotes input feature dimension. We choose the sparse support ${\mathcal{I}}$ uniformly at random up to the desired sparsity level $\delta$. Furthermore, in order to balance the contribution of low-rank factor and sparse factor, we follow LoRA [21] to scale the low-rank factors by $\alpha/r$ where the balancing parameter $\alpha$ is a hyperparameter. This hyperparameter, along with the stepsize has a joint effect on the training speed of low-rank versus sparse factors.
@@ 5.1 Pretraining LLMs
Memory cost estimation. We compare the proposed SLTrain with the low-rank baseline models in terms of estimated memory consumption. Following [59], we compute memory estimates with bfloat16 format, where each floating point number occupies 2 bytes. We remark that SLTrain stores the indices with int64 format, which occupies 8 bytes per digit. The memory cost for a training algorithm consists of the parameter memory and optimizer state memory. The parameter memory refers to the memory occupied by storing parameters, and the optimizer state memory refers to the memory required to store the first and second-order moment statistics, e.g., in Adam. Table 2 reports the total estimated memory cost for each method. The detailed breakdown of memory estimation can be found in Appendix F.
Perplexity vs efficiency. In Table 2, we compare the performance of different methods in three aspects: perplexity score, parameter size, and memory cost. We observe that SLTrain performs comparatively as the full-rank training and GaLore [59] while achieving further reduction in parameter size and memory cost. In addition, SLTrain only adds a small parameter and memory overhead to the low-rank parameterization, yet significantly improves the perplexity score. Hence, learning the additional sparse factor indeed helps in strengthening the representation capacity of SLTrain. This intuition is also validated in Figure 10 (Appendix D), where we plot the singular value distribution for different weight matrices. Due to the presence of the sparse factor, we observe that the spectrum of the SLTrain weight matrices gets enhanced beyond $r=128$.
@@
Scaling to LLaMA 7B model. For pretraining the LlaMA 7B model, due to the resource constraints, we only compare SLTrain with GaLore, implemented with 8-bit Adam [9] on 4$\times$80G A100 GPUs, without per-layer weight updates nor gradient checkpointing. [^1] For full-rank model, 8-bit Adam throws out-of-memory error on 4$\times$80G A100 GPUs.. We directly use the training scripts of GaLore. [^2] The scripts are available at https://github.com/jiaweizzhao/GaLore In Table 4, we observe that SLTrain performs comparably to GaLore in terms of perplexity and throughput while achieving significant memory reduction of 26% per GPU device.
Inference memory and throughput. In Table 5.1, we compare the inference memory usage and throughput between SLTrain and the full-rank model across various model sizes, ranging from LLaMA 130M to 7B. A clear trade-off between memory and computational cost is observed. Specifically, as the model size increases, the percentage of memory savings becomes more pronounced, while the corresponding increase in computational cost is less significant. This underscores the growing advantage of SLTrain when employing larger models.
@@
Table 5: Inference memory and throughput comparison on a single 40G A100 GPU on a batch size of 128 for 130M, 350M, and a batch size of 32 for 1B, 7B. Compared to 1B model, higher memory for 350M model is due to the larger batch size.
130M 350M
Mem Tokens/s Mem Tokens/s
Full-Rank 8.09G 151360 11.06G 71324
SLTrain 7.95G 137058 10.44G 66616
(-1.73%) (-9.45%) (-5.61%) (-6.60%)
1B 7B
Mem Tokens/s Mem Tokens/s
Full-Rank 8.64G 18964 32.93G 9500
SLTrain 6.12G 17482 21.19G 8481
(-29.17%) (-7.81%) (-35.65%) (-10.73%)
@@
How do rank $r$ and sparsity $\delta$ affect performance? In Table 6, we validate the performance of training the Llama 60M and 130M models by varying the hyperparameters $r$ and $\delta$. We notice that, in general, when more parameters are added (corresponding to higher $r$ and $\delta$), the performance is better. However, this performance gain is also accompanied by an increase in memory footprint. In our initial experiments, we also tried the extreme setting with $r=0$, i.e., only the $S$ component was learned but with a higher $\delta$ value. This setting demonstrated a reasonably good performance. Further investigations in this direction are left for future work.
Table 6: Ablation comparison with low-rank and sparse parameterization along with change of rank $r$ and sparsity $\delta$. Validation perplexity ($\downarrow$) and parameter size and estimated memory cost in brackets.
@@
Table 7: Results training LLaMA 350M (with batch size=128 per GPU) and LLaMA 1B (with batch size=32 per GPU). Validation perplexity (PPL) ()($\downarrow$)( ↓ ), number of parameters in millions (Param) and actual max memory allocated per GPU in G (Mem).
350M 1B
PPL Param Mem PPL Param Mem
Full-Rank 18.80 368 59.34 15.56 1339 39.97
SLTrain ($\delta=0.03$) 19.42 194 (-47%) 57.90 (-2.4%) 16.14 646 (-52%) 33.77 (-15.5%)
SLTrain ($\delta=0.05$) 19.24 200 (-45%) 58.00 (-2.2%) 15.97 670 (-50%) 34.30 (-14.2%)
SLTrain ($\delta=0.1$) 18.72 215 (-42%) 58.25 (-1.8%) 15.59 730 (-45%) 35.36 (-11.5%)
Further comparisons to full-rank performance.
In this section, we evaluate the potential of SLTrain to achieve performance comparable to full-rank models by adjusting the sparsity ratio, $\delta$. As shown in Table 5.1, we increase $\delta$ from 0.03 to 0.1 for the LLaMA 350M and 1B models. Our results indicate that setting $\delta=0.1$ enables SLTrain to attain perplexity scores similar to those of full-rank models, while maintaining memory and parameter efficiency. Notably, at $\delta=0.1$, SLTrain reduces the parameter size by 42% for the 350M model and 45% for the 1B model. These results highlight the effectiveness of SLTrain in significantly reducing model size without sacrificing performance.